# Controlling Ensemble Variance in Diffusion Models: An Application for Reanalyses Downscaling

### v1

@@ 5.1 Quantifying variance
Figure 3: Plotting the global mean variance $\mu_{V}$ of ensemble diffusion for different number of steps against CERRA-EDA
@@
Here, $\mu_{\boldsymbol{V}}$ is a scalar value representing the global mean variance across all samples and pixels. By computing $\mu_{\boldsymbol{V}}$ for different diffusion steps, we observe that the global variance increases with the number of steps, as illustrated in Figure 3. Furthermore, we demonstrate that the variance control mechanism successfully aligns the global variance with that of CERRA-EDA, with the closest match achieved at $N=12$ diffusion steps.
@@ 5.2 Analyzing variance spatially
Here, $\boldsymbol{V}_{i}$ and $\boldsymbol{\mu}^{spatial}_{k}$ are understood to represent spatial fields, where the value of each pixel is computed independently. We compute the spatial mean-variance for the reference ensemble CERRA-EDA and for Ensemble Diffusion across a range of diffusion steps $N\in\{2,4,8,10,12,14\}$ for the year 2016. This allows us to visualize how the spatial distribution of variance evolves as the number of diffusion steps increases. The obtained results are reported in Figure 4. We observe substantial overlap between the variance distributions of CERRA-EDA and Ensemble Diffusion. This indicates that the diffusion model successfully captures the variability of the data by training solely on the CERRA dataset, without requiring additional information about the underlying uncertainty. Furthermore, we note that the degree of spatial overlap varies by season, with the highest overlap observed in the JFM and OND sets, while AMJ and JAS exhibit less overlap.
@@
2 steps 4 steps 8 steps 10 steps 12 steps 14 steps
[…]
@@ 5.3 Impact of $\Delta t$ on performance
2 steps 4 steps 8 steps 10 steps 12 steps 14 steps
[…]
Table 3: Table comparing the MSE performance of the diffusion model at different numbers of reverse diffusion steps. The MSE is calculated between the mean of the 10 generated ensembles and the CERRA-EDA control run. The results are reported 3-monthly for the year 2016.
@@ 5.4 Demonstrating capabilities on CARRA
Both CERRA and CARRA, introduced in Section 2 are built upon the HARMONIE-AROME weather prediction system [50], adapted for reanalysis purposes. This common foundation ensures that both systems share a core methodology for atmospheric modeling and data assimilation, making them methodologically similar despite their focus on different regions.
@@ 6 Conclusions
[…]
@@
Since our diffusion model is trained solely on the deterministic ERA5-to-CERRA task, it is unaffected by factors such as measurement quality or other previously assessed uncertainties. As such, our approach can serve as a validation tool for existing ensembles, with discrepancies providing insights into the neural model’s understanding of the underlying physical system or serving as validation for uncertainty estimates obtained through other methods. In our experimental setting, the spatial variance reported in Figure 4 shows a clear overlap between distributions. However, during the summer months in the North African region, discrepancies between CERRA-EDA and our Ensemble Diffusion appear. This may indicate that our model struggles to capture uncertainty in this region, or that CERRA-EDA itself may be affected by data scarcity [6].
@@
In conclusion, our methodology provides a practical tool for tuning ensemble diffusion models to produce the correct variance distribution. We anticipate that future implementations will consider the number of diffusion steps as a crucial tuning parameter. Furthermore, our theoretical framework opens promising directions for studying the statistical properties of ensemble members in relation to the structure of diffusion models, encouraging further research.
@@ 7 Acknowledgements
The code and downscaled ensemble members relative to the presented work is archived at the following GitHub https://github.com/fmerizzi/ERA5-to-CERRA-via-Diffusion-Models/. We also provide the full list of the environmental libraries and respective versions.
@@ Appendix A Proofs of main results
A.1 Properties of element-wise variance
At first, we need to show that element-wise variance $\boldsymbol{v}(\boldsymbol{x}_{t})$, defined in (10) inheritates similar properties as the full variance, and in particular (P1) and (P2). This is not hard to prove, as:
$\boldsymbol{v}(a\boldsymbol{x}_{t})_{i}=Var(a\boldsymbol{x}_{t})_{i,i}=a^{2}Var(\boldsymbol{x}_{t})_{i,i}=a^{2}\boldsymbol{v}(\boldsymbol{x}_{t})_{i},$
and:
$\boldsymbol{v}(\boldsymbol{x}_{t}+\boldsymbol{y}_{t})_{i}$ $=Var(\boldsymbol{x}_{t}+\boldsymbol{y}_{t})_{i,i}=Var(\boldsymbol{x}_{t})_{i,i}+Var(\boldsymbol{y}_{t})_{i,i}+2Cov(\boldsymbol{x}_{t},\boldsymbol{y}_{t})_{i,i}$
$=\boldsymbol{v}(\boldsymbol{x}_{t})_{i}+\boldsymbol{v}(\boldsymbol{y}_{t})_{i}+2\boldsymbol{v}(\boldsymbol{x}_{t},\boldsymbol{y}_{t})_{i},$
which proves the required properties.
A.2 Proof of Proposition 3.1
The essence of our theoretical results lies on the recursive formula introduced in Proposition 3.1, which follows by properties (P1) and (P2) applied to the diffusion process (4). In particular note that, by (4),
$\boldsymbol{v}(\boldsymbol{x}_{t})=\boldsymbol{v}\left(\frac{\alpha_{t}}{\alpha_{t-\Delta t}}\boldsymbol{x}_{t-\Delta t}+c_{t-\Delta t}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})\right).$
Then, applying (P1) and (P2),
$\boldsymbol{v}(\boldsymbol{x}_{t})=\frac{\alpha_{t}}{\alpha_{t-\Delta t}}\boldsymbol{v}(\boldsymbol{x}_{t-\Delta t})+$ $c_{t-\Delta t}^{2}\boldsymbol{v}(\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t}))$
$+2\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}c_{t-\Delta t}\boldsymbol{v}(\boldsymbol{x}_{t-\Delta t},\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})).$
Since by (2), $\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})\approx\boldsymbol{\epsilon}_{t-\Delta t}$, then $\boldsymbol{v}(\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t}))\approx\boldsymbol{v}(\boldsymbol{\epsilon}_{t-\Delta t})=\boldsymbol{1}$, which results in:
$\boldsymbol{v}(\boldsymbol{x}_{t})=\frac{\alpha_{t}}{\alpha_{t-\Delta t}}\boldsymbol{v}(\boldsymbol{x}_{t-\Delta t})+c_{t-\Delta t}^{2}\boldsymbol{1}+2\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}c_{t-\Delta t}\boldsymbol{v}(\boldsymbol{x}_{t-\Delta t},\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})).$
To simplify this expression, recall the definition of the covariance:
$\boldsymbol{v}(\boldsymbol{x}_{t-\Delta t},\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t}))=\mathbb{E}\left[(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t})(\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})-\mathbb{E}[\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})])^{T}\right],$
where we denoted $\boldsymbol{m}_{t-\Delta t}:=\mathbb{E}\left[\boldsymbol{x}_{t-\Delta t}\right]$, and consider the first order Taylor decomposition of $\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})$ around $\boldsymbol{m}_{t-\Delta t}$:
$\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})\approx\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})+\nabla_{\boldsymbol{m}_{t-\Delta t}}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})\left(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t}\right).$

### v2

@@ 5.1 Quantifying variance
Figure 3: Plotting the global mean variance $\mu_{V}$ of ensemble diffusion for different number of steps against CERRA-EDA.
@@
Here, $\mu_{\boldsymbol{V}}$ is a scalar value representing the global mean variance across all samples and pixels. By computing $\mu_{\boldsymbol{V}}$ for different diffusion steps, we observe that the global variance increases with the number of steps, as illustrated in Figure 3. Furthermore, we demonstrate that the variance control mechanism successfully aligns the global variance with that of CERRA-EDA, with the closest match achieved at $N=12$ diffusion steps. Interestingly, the relationship between the number of steps and ensemble variance may have been overlooked in previous research, as it is common practice to run diffusion models with a high number of steps [ling2024diffusion, price2023gencast, bassetti2024diffesm]. In such cases, the overall variance would have already converged to a fixed value, making it unaffected by changes in the number of reverse diffusion steps.
@@ 5.2 Analyzing variance spatially
Here, $\boldsymbol{V}_{i}$ and $\boldsymbol{\mu}^{spatial}_{k}$ are understood to represent spatial fields, where the value of each pixel is computed independently. We compute the spatial mean-variance for the reference ensemble CERRA-EDA and for Ensemble Diffusion across a range of diffusion steps $N\in\{2,4,6,8,10,12,14\}$ for the year 2016. This allows us to visualize how the spatial distribution of variance evolves as the number of diffusion steps increases. The obtained results are reported in Figure 4. We observe substantial overlap between the variance distributions of CERRA-EDA and Ensemble Diffusion. This indicates that the diffusion model successfully captures the variability of the data by training solely on the CERRA dataset, without requiring additional information about the underlying uncertainty. Furthermore, we note that the degree of spatial overlap varies by season, with the highest overlap observed in the JFM and OND sets, while AMJ and JAS exhibit less overlap.
@@
2 steps 4 steps 6 steps 8 steps 10 steps 12 steps 14 steps
yearly $1.88\text{e-01}$ $1.45\text{e-01}$ $1.41\text{e-01}$ $\underline{1.40\text{e-01}}$ $1.43\text{e-01}$ $1.45\text{e-01}$ $1.49\text{e-01}$
JFM $1.14\text{e-01}$ $\underline{1.02\text{e-01}}$ $04\text{e-01} 1.1$$0\text{e-01} 1.1$$19\text{e-01}$ $1.27\text{e-01}$ $1.32\text{e-01}$
AMJ $2.15\text{e-01}$ $1.62\text{e-01}$ $1.42\text{e-01}$ $1.37\text{e-01}$ $1.34\text{e-01}$ $\underline{1.33\text{e-01}}$ $1.34\text{e-01}$
JAS $2.22\text{e-01}$ $1.74\text{e-01}$ $1.58\text{e-01}$ $1.50\text{e-01}$ $1.47\text{e-01}$ $\underline{1.46\text{e-01}}$ $1.47\text{e-01}$
@@
Figure 5: Evolution of the mean variance discrepancy (MVD) scores with the number of diffusion steps. The plot visualizes the values reported in Table 2 for the yearly and three-monthly periods (JFM, AMJ, JAS, OND) during 2016. Lower values indicate a closer match between the spatial variance distributions of Ensemble Diffusion and CERRA-EDA. The yearly curve (orange) highlights the global trend, while the seasonal curves illustrate the differing optimal step counts across the annual cycle.
@@ 5.3 Impact of $\Delta t$ on performance
For what concerns execution efficiency, the number of diffusion steps is expected to scale linearly with the total inference time. Under an inference setup analogous to the training configuration described in Section 4.2 (three NVIDIA A100 GPUs) the time required to generate 10 ensemble members for the year 2016 at a 6-hour temporal resolution ranges from approximately 110 minutes for 2 diffusion steps to 746 minutes for 14 steps. The relationship between inference time and the number of diffusion steps is illustrated in Figure 6.
Figure 6: Inference time of the ensemble diffusion model as a function of the number of diffusion steps. The results show an approximately linear relationship, confirming that computational cost scales proportionally with the number of reverse diffusion iterations.
@@
2 steps 4 steps 6 steps 8 steps 10 steps 12 steps 14 steps
yearly $1.67\text{e-03}$ $1.62\text{e-03}$ $1.60\text{e-03}$ $1.59\text{e-03}$ $1.58\text{e-03}$ $1.58\text{e-03}$ $\underline{1.57\text{e-03}}$
JFM $1.98\text{e-03}$ $1.92\text{e-03}$ $1.90\text{e-03}$ $1.88\text{e-03}$ $1.87\text{e-03}$ $1.86\text{e-03}$ $\underline{1.86\text{e-03}}$
AMJ $1.52\text{e-03}$ $1.48\text{e-03}$ $1.47\text{e-03}$ $1.46\text{e-03}$ $1.45\text{e-03}$ $1.45\text{e-03}$ $\underline{1.44\text{e-03}}$
JAS $1.39\text{e-03}$ $1.36\text{e-03}$ $1.34\text{e-03}$ $1.33\text{e-03}$ $1.33\text{e-03}$ $1.33\text{e-03}$ $\underline{1.32\text{e-03}}$
OND $1.79\text{e-03}$ $1.73\text{e-03}$ $1.70\text{e-03}$ $1.68\text{e-03}$ $1.68\text{e-03}$ $\underline{1.67\text{e-03}}$ $\underline{1.67\text{e-03}}$
Table 3: Table comparing the MSE performance of the diffusion model at different numbers of reverse diffusion steps. The MSE is calculated between the mean of the 10 generated ensembles and the CERRA-EDA control run. The results are reported 3-monthly for the year 2016, along with the yearly mean across all seasons.
Figure 7: Mean squared error (MSE) of the ensemble mean versus the CERRA-EDA control run for different diffusion steps. Results for each three-monthly period (JFM, AMJ, JAS, OND) in 2016 show stable performance with slightly lower errors at higher step counts.
@@ 5.4 Demonstrating capabilities on CARRA
Both CERRA and CARRA, introduced in Section 2 are built upon the HARMONIE-AROME weather prediction system [bengtsson2017harmonie], adapted for reanalysis purposes. This common foundation ensures that both systems share a core methodology for atmospheric modeling and data assimilation, making them methodologically similar despite their focus on different regions.
@@
5.5 Comparison with EMOS
Our method provides a direct ensemble comparison with CERRA-EDA, the numerically computed ensemble of CERRA, which can be naturally regarded as the optimal ground truth for evaluation. To ensure a comprehensive evaluation, our method can be positioned alongside other ensemble-generation approaches, including well-established statistical post-processing techniques such as EMOS. Among neural stochastic architectures, methods based on GANs have been proposed for related tasks, yet their training and tuning requirements make them impractical in our setting, as discussed in Section 1.1. For this reason, our comparison focuses on EMOS.
EMOS (Ensemble Model Output Statistics) is a statistical post-processing approach widely used to calibrate ensemble and deterministic forecasts [gneiting2005calibrated]. Originally developed as a calibration framework for correcting bias and spread in ensemble systems, EMOS has become one of the most adopted post-processing techniques in numerical weather prediction and downscaling contexts [thorarinsdottir2010probabilistic, baran2016mixture, szabo2023parametric]. Beyond calibration, EMOS can also be applied to generate probabilistic ensembles directly from deterministic forecasts by modeling the predictive distribution of the target variable as a function of the forecast itself. This formulation allows EMOS to correct systematic biases and rescale forecast variability, producing calibrated ensemble predictions that are statistically consistent with the observed data.
In our setting, EMOS is applied to the deterministic outputs of the reference U-Net to generate probabilistic ensembles. We employ a deterministic version of our downscaling model, a high-performing residual U-Net, which achieves performance comparable to the diffusion model in terms of MSE for the testing years, as reported in Table 1. This deterministic U-Net provides an ideal baseline for comparison, and our experiment focuses on generating ensemble members from the U-Net using EMOS and evaluating them against those obtained from the diffusion-based ensembles.
The objective of this experiment is to evaluate whether EMOS, when fitted to the relationship between the deterministic forecasts and the ground truth, can reproduce the ensemble variability observed in CERRA-EDA. In practice, EMOS models the conditional distribution of wind speed as the truncated normal distribution, $Y\sim\mathcal{TN}_{0}(\mu=a+bX,\sigma)$, where the coefficients $a$ and $b$ correct systematic bias and amplitude scaling, and $\sigma$ controls the ensemble spread. The deterministic U-Net, trained on the 1985–2010 period (consistent with the diffusion model), was used as the forecast source. The year 2015 was used to fit the EMOS parameters against the CERRA ground truth, while 2016 was reserved for evaluation against the ten-member CERRA-EDA ensemble. All datasets were normalized to the $[0,1]$ range using the CERRA training statistics. EMOS fitting was performed after bilinearly downscaling the raw U-Net outputs to $565\times 565$ pixels to match the spatial resolution of CERRA-EDA. The evaluation was carried out at a $6$-hourly frequency and $565\times 565$ spatial resolution, comparing the EMOS-generated ensemble with the reference ensemble dataset.
Total Ensemble Variance Comparison
2 steps $0.216$
4 steps $0.298$
6 steps $0.338$
8 steps $0.357$
10 steps $0.375$
12 steps $0.388$
14 steps $0.397$
EMOS $0.392$
Local EMOS $0.392$
CERRA-EDA (Ref.) $\mathbf{0.392}$
Table 4: Total ensemble variance ($\mu_{V}$) for Ensemble Diffusion at increasing numbers of reverse diffusion steps, compared with EMOS, Local EMOS, and the reference CERRA-EDA variance. Variance increases monotonically with diffusion steps and converges toward the reference.
Two EMOS configurations were tested. In the standard EMOS, a single set of parameters $(a,b,\sigma)$ was fitted over the entire 2015 dataset, yielding $a=-7.1\times 10^{-4}$, $b=1.015$, and $\sigma=0.0168$. The spread was subsequently calibrated to match the global mean variance of CERRA-EDA via $\sigma_{\mathrm{new}}=\sigma\sqrt{\mu V_{\mathrm{target}}/\mu V_{\mathrm{current}}}$, resulting in $\sigma_{\mathrm{new}}=0.0122$. In the local EMOS, a separate regression was fitted for each grid point, $Y_{t}(i,j)=a_{i,j}+b_{i,j}X_{t}(i,j)+\varepsilon_{t}(i,j)$, producing spatial maps of $a_{i,j}$, $b_{i,j}$, and $\sigma_{i,j}$. To prevent over-dispersion, the per-pixel spread was globally scaled using the same variance calibration factor, $k=\sqrt{\mu V_{\mathrm{target}}/\mu V_{\mathrm{current}}}\approx 0.60$.
MVD Scores for Ensemble Diffusion (8 steps) vs EMOS
8 steps (Ours) EMOS Local EMOS
Yearly $\underline{1.40\text{e-01}}$ $1.60\text{e-01}$ $1.59\text{e-01}$
JFM $\underline{1.11\text{e-01}}$ $1.65\text{e-01}$ $1.59\text{e-01}$
AMJ $\underline{1.37\text{e-01}}$ $1.51\text{e-01}$ $1.52\text{e-01}$
JAS $1.51\text{e-01}$ $\underline{1.40\text{e-01}}$ $1.44\text{e-01}$
OND $\underline{1.17\text{e-01}}$ $1.43\text{e-01}$ $1.40\text{e-01}$
Table 5: Mean variance discrepancy (MVD) scores computed between the spatial mean variance maps of the 8-step Ensemble Diffusion model, EMOS, and Local EMOS for each three-monthly period (JFM, AMJ, JAS, OND) and the full year 2016. Lower scores indicate higher similarity between the variance distributions of the compared ensembles.
Naturally, both EMOS variants successfully reproduce the total ensemble variance of CERRA-EDA, as they are directly calibrated to match this value, as reported in Table 4. The more interesting question is whether they can approximate the spatial distribution of the variance. For the standard EMOS approach, which relies on a single value of $\sigma$, this is inherently impossible. However, the local EMOS formulation, which estimates parameters for each grid point, may be able to capture some spatial structures. Having computed spatial variance maps analogous to those obtained for the diffusion ensembles, the local EMOS results were added to Figure 4. As shown in the figure, EMOS identifies a few regions of high variability, but these are limited in extent and fail to represent the complex spatial patterns of variance clearly captured by both CERRA-EDA and the diffusion model.
The mean-variance discrepancy (MVD) scores computed between EMOS, local EMOS, and the best-performing diffusion configuration (8 steps), reported in Table 5, confirm this observation. In the yearly comparison, the diffusion model clearly outperforms both EMOS configurations, and the same holds for three of the four seasonal partitions. Interestingly, however, during the July–August–September period, the EMOS configuration slightly outperforms diffusion. This finding is noteworthy, given that EMOS is a purely statistical method that does not explicitly model spatial variability. This outcome likely reflects a relative reduction in the diffusion model’s ability to capture uncertainty during the summer months, rather than a genuine advantage of EMOS, suggesting that ensemble diffusion may struggle to represent variance under the distinct dynamical conditions of this season.
5.6 Limitations
While the proposed framework demonstrates that ensemble diffusion models can effectively reproduce the statistical variability of regional wind fields, several aspects merit further exploration. First, the analysis has been limited to wind speed. Although preliminary tests on other variables such as temperature and precipitation yielded comparable results, suggesting that the core principles of our approach (variance calibration and spatial consistency evaluation) may generalize beyond wind-related fields, a comprehensive multi-variable analysis would provide valuable insight into the universality and robustness of the proposed methodology, including any variable-specific adjustments that may be required. The validation of ensemble variance relied on CERRA-EDA as the reference dataset. While this ensemble offers a consistent and physically grounded benchmark, it operates at a reduced spatial and temporal resolution and thus may not capture the full spectrum of uncertainty present in observations. Future studies could strengthen the evaluation by incorporating in-situ measurements, alternative reanalysis ensembles, or observational uncertainty estimates. Geographically, the experiments focused on the European CERRA domain with an extension to CARRA-EAST. Although these regions encompass diverse climatological regimes, confirming the model’s robustness across other continents and observational contexts (particularly areas with different dynamics or sparser data) would further consolidate the method’s general applicability. Finally, the evaluation primarily targeted variance alignment through global metrics such as mean variance deviation (MVD). While this provides a first-order measure of ensemble consistency, complementary diagnostics, including spread–skill relationships, reliability diagrams, and extreme-event statistics, could offer a more comprehensive assessment of ensemble quality. Overall, these limitations delineate promising directions for future work. Extending the framework to additional variables, reference datasets, and regions, as well as enriching the evaluation protocols, will further clarify the potential of diffusion-based ensembles as physically interpretable and controllable tools for climate downscaling.
@@ 6 Conclusions
In this work, we applied an ensemble diffusion model to the full-domain ERA5-to-CERRA downscaling task and discovered a relationship between the variance exhibited by the ensemble diffusion and the number of reverse diffusion steps, giving theoretical proof of this relation. Notably, we demonstrated that the selection of the reverse diffusion step size $\Delta t$ not only impacts the time efficiency of the generative process but also serves as a crucial mechanism for controlling the variance of the sampled data, a key aspect when using these models for generating ensembles. We have explored architectural improvements needed to apply ensemble diffusion models over large domains, including the use of a scaling factor to reduce the signal rate in the noise scheduler, reducing the effect that higher resolution have on denoising difficulty. Our experimental results indicate that adjusting the number of diffusion steps enables the model to operate across different ranges of ensemble variance. We found that the control over variance gained by varying the number of diffusion steps is sufficient to match the global mean variance exhibited by the available ensemble reference dataset, CERRA-EDA. We observed that increasing the number of diffusion steps leads to convergence toward a fixed variance value. Although this converged global mean variance closely matches the one exhibited by CERRA-EDA, our MVD score, which takes into account the spatial distribution of the variance, revealed that the optimal number of steps is 8, a value far from the convergence point. This demonstrates that selecting a “sufficiently high” number of steps is not necessarily an optimal strategy for tuning an ensemble diffusion model, highlighting the importance of calibration to obtain optimal results.
To further contextualize the effectiveness of our approach, we have also introduced an Ensemble Model Output Statistics (EMOS) baseline as a comparison method. EMOS, a widely adopted statistical post-processing technique for ensemble calibration, provided a valuable reference for assessing the seasonal variability and overall performance of our diffusion model. While EMOS successfully reproduced global variance levels, it lacked the spatial variance coherence achieved by the diffusion-based approach, underscoring the advantages of learning-based generative methods for ensemble generation.
In our work we have showcased the capability of generating ensemble members efficiently at high temporal and spatial resolutions. Specifically, our ensemble reference CERRA-EDA has a considerable reduction in spatial resolution (from $1069\times 1069$ to $565\times 565$) and temporal resolution (from $3h$ to $6h$). We reveal that our approach is able to generate ensemble members at full CERRA temporal and spatial resolution, while maintaining high efficiency, essentially running on workstation hardware. Considering that our model is conditioned only by ERA5 low-resolution inputs, which happens to be available at hourly intervals, we can additionally increase the ensemble temporal resolution to $1h$. Additionally, we have extended our ensemble diffusion methodology from CERRA to CARRA-EAST, demonstrating its capability to generate ensemble members for datasets lacking ensemble information, using a similar training setup and confirming its effectiveness in reproducing spatial variance patterns through diffusion step tuning.
@@
Since our diffusion model is trained solely on the deterministic ERA5-to-CERRA task, it is unaffected by factors such as measurement quality or other previously assessed uncertainties. As such, our approach can serve as a validation tool for existing ensembles, with discrepancies providing insights into the neural model’s understanding of the underlying physical system or serving as validation for uncertainty estimates obtained through other methods.
In our experimental setting, the spatial variance reported in Figure 4 shows a clear overlap between distributions. However, during the summer months in the North African region, discrepancies between CERRA-EDA and our Ensemble Diffusion appear. This may stem from physical factors limiting the model’s ability to learn regional variance dynamics, particularly under intense surface heating, shallow convective mixing, and heterogeneous land-surface conditions typical of the summer season. Alternatively, it may reflect inflated variance in CERRA-EDA itself, driven by data scarcity and uncertainty inflation in areas with sparse observations. Although a full investigation of these mechanisms lies beyond the scope of this work, we note that a season-specific tuning of the number of reverse diffusion steps mitigates this issue.
@@
In conclusion, our methodology provides a practical tool for tuning ensemble diffusion models to produce the correct variance distribution. We believe that future implementations should consider the number of diffusion steps as a crucial tuning parameter. Furthermore, our theoretical framework opens promising directions for studying the statistical properties of ensemble members in relation to the structure of diffusion models, encouraging further research.
@@ 7 Acknowledgements
Code availability removed to ensure blindness.
@@ Appendix A Proofs of main results
A.1 Full proof of Theorem 3.2 (scalar case)
For completeness, we report here the full derivation of Theorem 3.2, whose main steps were only sketched in the paper body. The proof below includes all intermediate Taylor expansions and variance–recursion arguments.
Proof.
Let $\alpha(t)$ be the continuous extension of $\{\alpha_{t}\}_{t=1}^{T}$. The first-order Taylor decomposition of $\sqrt{\alpha(t)}$ around $t-\Delta t$ shows:
$\sqrt{\alpha(t)}$ $=\sqrt{\alpha(t-\Delta t)}-\left(\sqrt{\alpha(t-\Delta t)}\right)^{\prime}\Delta t+O(\Delta t^{2})$
$=\sqrt{\alpha(t-\Delta t)}-\frac{\alpha^{\prime}(t-\Delta t)}{2\sqrt{\alpha(t-\Delta t)}}\,\Delta t+O(\Delta t^{2}),$
and similarly, for $\sqrt{1-\alpha(t)}$:
$\sqrt{1-\alpha(t)}$ $=\sqrt{1-\alpha(t-\Delta t)}-\left(\sqrt{1-\alpha(t-\Delta t)}\right)^{\prime}\Delta t+O(\Delta t^{2})$
$=\sqrt{1-\alpha(t-\Delta t)}+\frac{\alpha^{\prime}(t-\Delta t)}{2\sqrt{1-\alpha(t-\Delta t)}}\,\Delta t+O(\Delta t^{2}).$